# On the Price of Decentralization in Decentralized Detection

### v1

@@ IV-A Estimating the Optimal Geometric Weights in a Decentralized Way
Theorem 2 and Theorem 3 state that each node obtains the optimal error exponent as long as each node $i$ has access to $\pi_{i}$, the $i$-th entry of the stationary distribution of the Markov chain whose transition matrix is the weight matrix $W$. For each node to learn the local stationary distribution, some additional effort is needed in the decentralized setting. The most naive way is to request a center that knows the entire weight matrix $W$ to calculate the stationary distribution and disseminate the corresponding information to each node. However, such a centralized method is not desirable from the perspective of decentralization.
@@
$\leq$ $\;\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\pi}_{i}^{(0)}\left|[W^{T}]_{ij}-\pi_{i}\right|\leq\left(\sum_{i=1}^{n}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}}\right)\left(\sum_{i=1}^{n}\pi_{i}\left(\hat{\pi}_{i}^{(0)}\right)^{2}\right)\rho^{T}.$
@@
IV-B Combining the Learning Rule with the Estimation on the Optimal Geometric Weights
In the previous section, we provide a simple method for estimating the stationary distribution in a decentralized manner. Though the estimation error vanishes exponentially, a first-order loss remains due to the difference between the estimation and the true stationary distribution. However, we can keep estimating the stationary distribution while executing the social learning rule, and the combined learning rule becomes the following.
@@ IV-B Combining the Learning Rule with the Estimation on the Optimal Geometric Weights
In the first part, we estimate the stationary distribution for $T_{\mathsf{E}}$ rounds, and in the second part, we adopt our learning rule together with estimating the stationary distribution iteratively. The combined learning rule obtains the optimal error exponent, which could be shown with a proof similar to the one of Theorem 2.
@@ V-A Impact of Network Imbalance
In Section II-D, we show how network imbalance can impact the error exponent through Theorem 1 and Figure 1. In the following we provide a simulation for the impact of network imbalance to support our statement regarding the non-asymptotic performance.
In Figure 5, random scale-free networks with 30 nodes in each are sampled. Each node in the sampled network distributes its relative confidence uniformly to its neighbors to form the weight matrix, and the corresponding stationary distribution is evaluated. Each sampled network is tagged with its quantity of imbalance, which is the total variation between its corresponding stationary distribution and the uniform one, and clustered into five groups. The simulation results for these groups of networks are gathered in the five subplots accordingly. For example, the leftmost subplot gathers the results of 30 sampled networks with quantities of imbalance falling within 0.25 to 0.35. Notice that we use the total variation between the stationary distribution and the uniform one to quantify the network imbalance, however, one can use other distance measurements such as vector 2-norm.
@@ V-B Compensating the Network Imbalance
    Our learning rule with each node knowing the stationary distribution in prior.
@@ V-C The Effect of Initial and Ongoing Estimations
In Section IV-B, we show that we can estimate the stationary distribution for several rounds before each node starts drawing observations. Furthermore, we can choose whether to keep estimating the stationary distribution. In the following, we demonstrate the effect of such initial and ongoing estimations on the stationary distribution.
@@
The above results show that we obtain improvements if each node either executes a single round of initial estimation or just keeps estimating the stationary distribution while drawing the observations. However, our theoretical results show that each node obtains a loss in the error exponent if it does not keep estimating the stationary distribution. That is, we can expect that the dashed lines would eventually divert further from the light blue line (one with the prior knowledge on the stationary distribution) and be outperformed by other solid lines.
@@ VI-A Removing the Assumptions on the Network
We assume that the Markov chain governing the information consensus is both irreducible and aperiodic in Assumption 2. In this section, we show how our learning rule performs without these assumptions.
Irreducible.
@@ VI-B Multiple Hypothesis Testing
$P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}<\gamma^{*}_{i}(\theta_{k},\theta_{l})\right\}<\epsilon\leq P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}\leq\gamma^{*}_{i}(\theta_{k},\theta_{l})\right\}.$
@@
$P_{k}\left\{\hat{\theta}_{i}=\theta_{l}\right\}\leq$ $\;P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}<\gamma^{*}_{i}(\theta_{l},\theta_{k})\right\}\doteq e^{-tD_{\mathrm{KL}}(P_{l}\|P_{k})}$
@@ VII Conclusion
We study the price of decentralization in distributed hypothesis testing. The original learning rule introduced in Section II-B obtains a tilted statistic compared to the centralized one and thus leads to a sub-optimal error exponent. The sub-optimal result comes from the network imbalance, and we compensate for the imbalance by properly choosing the additional geometric weights introduced in our modified learning rule to achieve the optimal error exponent. Furthermore, we look for the higher-order asymptotics of the type-II error probability and the Bayes risk. We reveal an upper bound on the price of decentralization as a constant term in the exponent of the error probability, where the extra term depends on the connectivity of the underlying network and the network imbalance.
@@
There are also some future directions. First, our first-order results are optimal, while our analysis on the higher-order terms turns out to be upper bounds on the error probability. Whether a constant punishment is inevitable (except for a complete graph with uniform weights) remains unclear, and a non-trivial lower bound on the error probability is expected to answer the question. Secondly, in Section VI-B, we utilize the re-weighting method for the multiple hypothesis testing problem. Furthermore, we can consider the decentralized estimation problems where exchanging beliefs would not be viable. Other first-order methods such as decentralized gradient descent may also be considered.
@@ A-A Proof of Theorem 1
$\alpha_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})=P_{0}\left\{\phi_{i}^{(t)}(\ell_{i}^{(t)})=1\right\}=P_{0}\left\{\ell_{i}^{(t)}>\gamma_{i}^{(t)}\right\}+\eta_{i}^{(t)}P_{0}\left\{\ell_{i}^{(t)}=\gamma_{i}^{(t)}\right\}\leq P_{0}\left\{\ell_{i}^{(t)}\geq\gamma_{i}^{(t)}\right\},$
@@
$\beta_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})=P_{1}\left\{\phi_{i}^{(t)}(\ell_{i}^{(t)})=1\right\}\leq P_{1}\left\{\ell_{i}^{(t)}\leq\gamma_{i}^{(t)}\right\},$ (1)
@@
$\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})\leq$ $\;P_{1}\left\{\sum_{j=1}^{n}\sum_{\tau=1}^{t}[W^{\tau}]_{ij}r_{j}\log\frac{P_{j,1}(X_{j}^{(t-\tau+1)})}{P_{j,0}(X_{j}^{(t-\tau+1)})}\leq\gamma_{i}^{(t)}\right\}$
@@
To bound the probability in (2), we introduce a large deviation analysis.
For simplicity, let the random vector $Y_{i}^{(t)}$ be
@@
$Y\triangleq\left[Y_{1}\;\dots\;Y_{n}\right]^{\mathrm{T}},$
$Y_{j}\triangleq\lim_{t\rightarrow\infty}Y_{ij}^{(t)}\sim\pi_{j}r_{j}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}.$
@@
$Z_{i}^{(t)}\triangleq\left[Z_{i1}^{(t)}\;\dots\;Z_{in}^{(t)}\right]^{\mathrm{T}},$
$Z_{ij}^{(t)}\triangleq\frac{1}{t}\sum_{\tau=1}^{t}Y_{ij}^{(\tau)}.$
@@
$\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})\leq P_{1}\left\{\sum_{j=1}^{n}\sum_{\tau=1}^{t}Y_{ij}^{(t)}\leq\gamma_{i}^{(t)}\right\}=P_{1}\left\{\sum_{j=1}^{n}Z_{ij}^{(t)}\leq\frac{1}{t}\gamma_{i}^{(t)}\right\}.$
@@
$\Lambda(\lambda)\triangleq$ $\;\lim_{t\rightarrow\infty}\frac{1}{t}\Lambda_{t}(t\lambda)=\lim_{t\rightarrow\infty}\frac{1}{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle t\lambda,Z_{i}^{(t)}\rangle}\right]=\lim_{t\rightarrow\infty}\frac{1}{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,\sum_{\tau=1}^{t}Y_{i}^{(\tau)}\rangle}\right]$
$=$ $\;\lim_{t\rightarrow\infty}\frac{1}{t}\log\left(\prod_{\tau=1}^{t}\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,Y_{i}^{(\tau)}\rangle}\right]\right)=\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{\tau=1}^{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,Y_{i}^{(\tau)}\rangle}\right],$ (3)
@@
$\gamma_{i}^{(t)}=t\left(\sum_{j=1}^{n}\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\right]+\delta\right),$
@@
=$\mathcal{F}=\left\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}\leq\gamma_{i}.\right\}$
@@
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})\geq$ $\;\liminf_{t\rightarrow\infty}-\frac{1}{t}\log P_{1}\left\{\sum_{j=1}^{n}Z_{ij}^{(t)}\leq\gamma_{i}\right\}\geq\inf_{x\in\mathcal{F}}\Lambda^{*}(x)$
@@
Since the logarithmic moment generating function $\Lambda(\lambda)$ is convex in $\lambda$, by minimax theorem, we exchange the order of the infimum and supremum in (4). In (5), observe that if $\lambda$ has any positive entry, the term $\langle\lambda,x\rangle$ easily goes to negative infinity, and thus the optimal $\lambda$ must fall in $\mathbb{R}_{\leq 0}^{n}$. Furthermore, if $\lambda$ is not orthogonal to the boundary of $\mathcal{F}$, that is, $x_{1}+\dots+x_{n}=\gamma_{i}$, we can always find an $x\in\mathcal{F}$ such that $\langle\lambda,x\rangle$ goes to negative infinity. Let $\boldsymbol{1}\in\mathbb{R}^{n}$ denote the vector with all entries being 1. Thus in (6) we simplify the optimization problem Instead of optimizing over $\lambda\in\mathbb{R}_{\leq 0}^{n}$. Now we can see that $x^{*}$ is optimal if and only if $x^{*}_{1}+\dots+x^{*}_{n}=\gamma_{i}$, thus we choose $x=\frac{\gamma_{i}}{n}\boldsymbol{1}$ and plug it in into (6). Plug the $\gamma_{i}$ we chose earlier in, and we have
$\sup_{\lambda\leq 0}\left\{\lambda\gamma_{i}-\log\mathsf{E}_{P_{1}}\left[e^{\lambda\sum_{j=1}^{n}Y_{j}}\right]\right\}=$ $\;\sup_{\lambda\leq 0}\left\{\begin{array}[]{l}\lambda\left(\sum_{j=1}^{n}\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\right]+\delta\right)\\ -\sum_{j=1}^{n}\log\mathsf{E}_{P_{1}}\left[e^{\lambda Y_{j}}\right]\end{array}\right\}$ (9)
@@
where since we can make $\delta$ arbitrarily close to zero, we can omit the first term in (15) and we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\beta^{(t)*}_{i}(r,\epsilon)\geq\sup_{\lambda\geq 0}\left\{\sum_{j=1}^{n}\lambda\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\log\mathsf{E}_{P_{1}}\left[\exp\left(\lambda\pi_{j}r_{j}\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\right]\right\}.$ (16)
The upper bound on the error exponent is obtained by the other part of the Gärtner-Ellis Theorem [10] with a similar technique.

### v2

@@ IV-A Estimating the Optimal Geometric Weights in a Decentralized Way
Theorem 2 and Theorem 3 state that the optimal error exponent can be attained at each node as long as each node $i$ has access to $\pi_{i}$, the $i$-th entry of the stationary distribution of the Markov chain whose transition matrix is the weight matrix $W$. For each node to learn the local stationary distribution, some additional effort is needed in the decentralized setting. The most naive way is to request a center that knows the entire weight matrix $W$ to calculate the stationary distribution and disseminate the corresponding information to each node. However, such a centralized method is not desirable from the perspective of decentralization.
@@
$\leq\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\pi}_{i}^{(0)}\left|[W^{T}]_{ij}-\pi_{i}\right|$
$\leq\left(\sum_{i=1}^{n}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}}\right)\left(\sum_{i=1}^{n}\pi_{i}\left(\hat{\pi}_{i}^{(0)}\right)^{2}\right)\rho^{T}.$
@@
IV-B Combining the Learning Rule with the Estimation of the Optimal Geometric Weights
In Section IV-A, we provide a simple method for estimating the stationary distribution in a decentralized manner. Though the estimation error vanishes exponentially, a first-order loss remains due to the difference between the estimation and the true stationary distribution. However, we can keep estimating the stationary distribution while executing the social learning rule, and the combined learning rule becomes the following.
@@ IV-B Combining the Learning Rule with the Estimation of the Optimal Geometric Weights
In the first part, we estimate the stationary distribution for $T_{\mathsf{E}}$ rounds, and in the second part, we adopt our learning rule together with estimating the stationary distribution iteratively. We conjecture that the combined learning rule achieves the optimal error exponent, which might be shown with a proof similar to the one of Theorem 2. In the next section, the performance of the proposed scheme is demonstrated through simulations.
@@ V-A Impact of Network Imbalance
We have shown that how network imbalance can impact the error exponent through Theorem 1 and Figure 1. In the following we provide a simulation for the impact of network imbalance to support our statement regarding the non-asymptotic performance.
In Figure 5, random scale-free networks with 30 nodes in each are sampled. Each node in the sampled network distributes its relative confidence uniformly to its neighbors to form the weight matrix, and the corresponding stationary distribution is evaluated. Each sampled network is tagged with its quantity of imbalance, which is the total variation between its corresponding stationary distribution and the uniform one, and clustered into five groups. The simulation results for these groups of networks are gathered in the five subplots accordingly. For example, the leftmost subplot gathers the results of 30 sampled networks with quantities of imbalance falling within 0.25 to 0.35. Notice that we use the total variation between the stationary distribution and the uniform one to quantify the network imbalance, however, one can use other distance measurements such as the vector 2-norm as in Figure 1.
@@ V-B Compensating the Network Imbalance
    Our learning rule with each node knowing the stationary distribution a priori.
@@ V-C The Effect of Initial and Ongoing Estimations
In the method proposed in Section IV-B, we estimate the stationary distribution for several rounds before each node starts drawing observations. Furthermore, we can choose whether to keep estimating the stationary distribution. In the following, we demonstrate the effect of such initial and ongoing estimations on the stationary distribution.
@@
The above results show that we obtain improvements if each node either executes a single round of initial estimation or just keeps estimating the stationary distribution while drawing the observations. However, it is anticipated that each node suffers a loss in the error exponent if it does not keep estimating the stationary distribution due to the estimation error of the optimal geometric weights. That is, we expect that the dashed lines would eventually divert further from the light blue line (one with the prior knowledge on the stationary distribution) and be outperformed by other solid lines.
@@ VI-A Removing the Assumptions on the Network
We assume that the Markov chain governing the information consensus is both irreducible and aperiodic in Assumption 2. In this section, we discuss how our learning rule performs without these assumptions.
Irreducible
@@ VI-B Multiple Hypothesis Testing
$P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}<\gamma^{*}_{i}(\theta_{k},\theta_{l})\right\}<\epsilon\quad\text{and}$
$P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}\leq\gamma^{*}_{i}(\theta_{k},\theta_{l})\right\}\geq\epsilon.$
@@
$P_{k}\left\{\hat{\theta}_{i}=\theta_{l}\right\}$ $\leq P_{k}\left\{\log\frac{b^{(t)}_{i}(\theta_{k})}{b^{(t)}_{i}(\theta_{l})}<\gamma^{*}_{i}(\theta_{l},\theta_{k})\right\}$
$\doteq e^{-tD_{\mathrm{KL}}(P_{l}\|P_{k})}$
@@ VII Conclusion
In this work, we study the price of decentralization in distributed hypothesis testing. The original learning rule introduced in Section II-B obtains a tilted statistic compared to the centralized one and thus leads to a sub-optimal error exponent. The sub-optimal result comes from the network imbalance, and we compensate for the imbalance by properly choosing the additional geometric weights introduced in our modified learning rule to achieve the optimal error exponent. Furthermore, we look for the higher-order asymptotics of the type-II error probability and the Bayes risk. We reveal an upper bound on the price of decentralization as a constant term in the exponent of the error probability, where the extra term depends on the connectivity of the underlying network and the network imbalance.
@@
Some directions are left as future work. First, our first-order results are optimal, while our analysis on the higher-order terms turns out to be upper bounds on the error probability. Whether a constant penalty is inevitable (except for a complete graph with uniform weights) remains unclear, and it is anticipated that a non-trivial lower bound on the error probability is needed to resolve the question. Secondly, despite the promising simulation results shown in Section V, a rigorous analysis on the probability of error while estimating the unknown geometric weight is lacking. Last but not least, taking the communication cost into account is an important next step towards a refined understanding of the price of decentralization in decentralized learning in practice.
@@ A-A Proof of Theorem 1
$\alpha_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})$ $=P_{0}\left\{\phi_{i}^{(t)}(\ell_{i}^{(t)})=1\right\}$
$=P_{0}\left\{\ell_{i}^{(t)}>\gamma_{i}^{(t)}\right\}+\eta_{i}^{(t)}P_{0}\left\{\ell_{i}^{(t)}=\gamma_{i}^{(t)}\right\}$
$\leq P_{0}\left\{\ell_{i}^{(t)}\geq\gamma_{i}^{(t)}\right\},$
@@
$\beta_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})$ $=P_{1}\left\{\phi_{i}^{(t)}(\ell_{i}^{(t)})=1\right\}$
$\leq P_{1}\left\{\ell_{i}^{(t)}\leq\gamma_{i}^{(t)}\right\},$ (2)
@@
$\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})$
$\leq P_{1}\left\{\sum_{j=1}^{n}\sum_{\tau=1}^{t}[W^{\tau}]_{ij}r_{j}\log\frac{P_{j,1}(X_{j}^{(t-\tau+1)})}{P_{j,0}(X_{j}^{(t-\tau+1)})}\leq\gamma_{i}^{(t)}\right\}$
@@
To bound the probability in (3), we introduce a large deviation analysis. For simplicity, let the random vector $Y_{i}^{(t)}$ be
@@
$Y\triangleq\left[Y_{1}\;\dots\;Y_{n}\right]^{\mathrm{T}},\ Y_{j}\triangleq\lim_{t\rightarrow\infty}Y_{ij}^{(t)}\sim\pi_{j}r_{j}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}.$
@@
$Z_{i}^{(t)}\triangleq\left[Z_{i1}^{(t)}\;\dots\;Z_{in}^{(t)}\right]^{\mathrm{T}},\ Z_{ij}^{(t)}\triangleq\frac{1}{t}\sum_{\tau=1}^{t}Y_{ij}^{(\tau)}.$
@@
$\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})$ $\leq P_{1}\left\{\sum_{j=1}^{n}\sum_{\tau=1}^{t}Y_{ij}^{(t)}\leq\gamma_{i}^{(t)}\right\}$
$=P_{1}\left\{\sum_{j=1}^{n}Z_{ij}^{(t)}\leq\frac{1}{t}\gamma_{i}^{(t)}\right\}.$
@@
$\Lambda(\lambda)$ $\triangleq\lim_{t\rightarrow\infty}\frac{1}{t}\Lambda_{t}(t\lambda)=\lim_{t\rightarrow\infty}\frac{1}{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle t\lambda,Z_{i}^{(t)}\rangle}\right]$
$=\lim_{t\rightarrow\infty}\frac{1}{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,\sum_{\tau=1}^{t}Y_{i}^{(\tau)}\rangle}\right]$
$=\lim_{t\rightarrow\infty}\frac{1}{t}\log\left(\prod_{\tau=1}^{t}\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,Y_{i}^{(\tau)}\rangle}\right]\right)$
$=\lim_{t\rightarrow\infty}\frac{1}{t}\sum_{\tau=1}^{t}\log\mathsf{E}_{P_{1}}\left[e^{\langle\lambda,Y_{i}^{(\tau)}\rangle}\right],$ (4)
@@
To this end, let us recall the Gärtner-Ellis Theorem from the large deviation theory as follows.
@@
$\gamma_{i}^{(t)}=t\left(\sum_{j=1}^{n}\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\right]+\delta\right)$
@@
=$\mathcal{F}=\left\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}\leq\gamma_{i}.\right\}.$
@@
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})$
$\geq\liminf_{t\rightarrow\infty}-\frac{1}{t}\log P_{1}\left\{\sum_{j=1}^{n}Z_{ij}^{(t)}\leq\gamma_{i}\right\}$
$\geq\inf_{x\in\mathcal{F}}\Lambda^{*}(x)$
@@
Here $\mathbb{R}_{\leq 0}$ denotes the set of non-positive real numbers. Since the logarithmic moment generating function $\Lambda(\lambda)$ is convex in $\lambda$, by the minimax theorem, we exchange the order of the infimum and supremum in (5). In (6), observe that if $\lambda$ has any positive entry, the term $\langle\lambda,x\rangle$ easily goes to negative infinity, and thus the optimal $\lambda$ must fall in $\mathbb{R}_{\leq 0}^{n}$. Furthermore, if $\lambda$ is not orthogonal to the boundary of $\mathcal{F}$, that is, $x_{1}+\dots+x_{n}=\gamma_{i}$, we can always find an $x\in\mathcal{F}$ such that $\langle\lambda,x\rangle$ goes to negative infinity. Let $\boldsymbol{1}\in\mathbb{R}^{n}$ denote the vector with all entries being 1. Thus in (7) we simplify the optimization problem Instead of optimizing over $\lambda\in\mathbb{R}_{\leq 0}^{n}$. Now we can see that $x^{*}$ is optimal if and only if $x^{*}_{1}+\dots+x^{*}_{n}=\gamma_{i}$, thus we choose $x=\frac{\gamma_{i}}{n}\boldsymbol{1}$ and plug it in into (7). Plug the $\gamma_{i}$ we chose earlier in, and we have
$\sup_{\lambda\leq 0}\left\{\lambda\gamma_{i}-\log\mathsf{E}_{P_{1}}\left[e^{\lambda\sum_{j=1}^{n}Y_{j}}\right]\right\}$
$=\sup_{\lambda\leq 0}\left\{\begin{array}[]{l}\lambda\left(\sum_{j=1}^{n}\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\right]+\delta\right)\\ -\sum_{j=1}^{n}\log\mathsf{E}_{P_{1}}\left[e^{\lambda Y_{j}}\right]\end{array}\right\}$ (10)
@@
Since we can make $\delta$ arbitrarily close to zero, we can omit the first term in (16) and we have
$\liminf_{t\rightarrow\infty}\frac{1}{t}\beta^{(t)*}_{i}(r,\epsilon)$
$\geq\sup_{\lambda\geq 0}\left\{\begin{array}[]{l}\sum_{j=1}^{n}\lambda\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]\\ -\log\mathsf{E}_{P_{1}}\left[\exp\left(\lambda\pi_{j}r_{j}\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\right]\end{array}\right\}.$ (19)
The upper bound on the error exponent is obtained by the other part of the Gärtner-Ellis Theorem [10] with a similar technique and hence omitted here.